John has $10,220 to invest. He invests part of his money in a CD at 5% interest and the rest in a savings account at 3.5%. If he earns $442.60 after on year, how much was invested in each account?
Answer by checkley77(12844)

.05x+.035(10,220-x)=442.60
.05x+357.70-.035=442.60
.015x=442.60-357.70
.015x=84.90
x=84.90/.015
x=$5,660 amount invested @ 5%.
10,220-5,660=4,560 amount invested @ 3.5%
Proof:
.05*5,660+.035*4,560=442.60
283+159.60=442.60
442.60=442.60